A car went 372 miles in 6 hours, traveling part of the time at 55 miles per hour and part of the time at 70 miles per hour. How long did the car travel at each speed?
Found 2 solutions by chiexpert, mananth:
Answer by chiexpert(48)

A car went 372 miles in 6 hours, traveling part of the time at 55 miles per hour and part of the time at 70 miles per hour. How long did the car travel at each speed?
0 solutions
D=RT
1ST -------D=55T
2ND----------372-D=70(6-T)
372-55T = 70(6) -70T
372-420=-70T+55T
48 = 15T
48/15=T 3.2 HOURS
3 HOURS AND 1/5 OF AN HOUR OR 12 MINUTES FOR THE FIRST 55 MPH
AND 3HOURS AND 48 MINUTES AT 70 MPH
According to my grid, "t" stands for the time spent on the first part of the trip, so my answer is "The CAR DROVE for 3 hours AND 12 MINUTES at 55 mph and 2 HOurs AND 48 MINUTES at 70 mph."

Answer by mananth(16949)

A car went 372 miles in 6 hours, traveling part of the time at 55 miles per hour and part of the time at 70 miles per hour. How long did the car travel at each speed?
..
let the car travel x miles at 55mph
the car will travel 372-x miles at 70 mph.
..
time at 55 mph = x/55
time at 70 mph = 372-x/70
..
x/55 + (372-x)/70 = 6
..
70x+55*372-55x / 70*55=6
15x+20460 =6*70*55
15x+20460=23100
add -20460
15x+20460-20460 = 23100-20460
15x=2640
divide by 15
x= 2640/15
x=176 miles at 55mph
372-x = 372-176= 196 miles at 70 mph